5. For an ideal Diesel cycle with the over-all value of k=1.33, rk=15, rc=2.1, P 1 = 14.2031 psia, find P 2 and

mean effective pressure.


1
Expert's answer
2021-04-23T08:25:16-0400

"P_1 =97.92 kPa"

"r_p = 2.1"


"P_2 = P_1\times (r_k)^{\gamma-1} = P_1\times (r_k)^{\gamma-1}"


"P_2 = 97.92 \times (15)^{1.33-1}=239.32kPa"


"P_m = \frac{P_1r_k(r_p-1)(r_k^{(\gamma-1) -1})}{(\gamma-1)(r_k-1)}"


"P_m = \frac{97.92 \times 15 (2.1-1)(15^{(1.33-1) }-1)}{(1.33-1)(15-1)} = 505 kPa"